# Does CLIP perceive art the same way we do?

### v1

@@
Abstract.
CLIP has emerged as a powerful multimodal model capable of connecting images and text through joint embeddings, but to what extent does it “see” the same way humans do—especially when interpreting artworks? In this paper, we investigate CLIP’s ability to extract high-level semantic and stylistic information from paintings, including both human-created and AI-generated imagery. We evaluate its perception across multiple dimensions—content, scene understanding, artistic style, historical period, and the presence of visual deformations or artifacts. By designing targeted probing tasks and comparing CLIP’s responses to human annotations and expert benchmarks, we explore its alignment with human perceptual and contextual understanding. Our findings reveal both strengths and limitations in CLIP’s visual representations, particularly in relation to aesthetic cues and artistic intent. We further discuss the implications of these insights for using CLIP as a guidance mechanism during generative processes, such as style transfer or prompt-based image synthesis. Our work highlights the need for deeper interpretability in multimodal systems, especially when applied to creative domains where nuance and subjectivity play a central role.
CLIP, multimodal models, painting analysis, generative guidance, vision-language alignment, computational art, visual perception
ccs: Computing methodologies Computer visionccs: Computing methodologies Scene understandingccs: Computing methodologies Neural networksccs: Computing methodologies Learning latent representationsccs: Computing methodologies Visual content-based indexing and retrievalccs: Information systems Multimedia content creationccs: Applied computing Fine arts
@@ Abstract.
Figure 1. 3D UMAP projection of image embeddings of National Gallery of Art Dataset extracted from the CLIP ViT-L/14 model.
[…]
Each point represents a painting from the National Gallery of Art dataset, colored by artistic style. The visualization reveals the semantic organization of the artworks in the latent space, where similar styles tend to cluster together.
1. Introduction
@@
In recent years, multimodal models have reshaped the landscape of machine perception and understanding, with CLIP (Contrastive Language–Image Pretraining) (Radford et al., 2021) standing out as one of the most influential. Trained on hundreds of millions of image–text pairs, CLIP has demonstrated remarkable capabilities across a broad range of tasks, including zero-shot classification (Wang et al., 2025; Li et al., 2025; Yang et al., 2024), image retrieval (Yang et al., 2025; Baldrati et al., 2024; Liu et al., 2024), re-identification (Li et al., 2023; Asperti et al., 2025b), prompt-based generation (Surapaneni and Bhagvati, 2024; Huang et al., 2023), and semantic search (Peng et al., 2024; Zhou et al., 2024). Its success has made it a foundational component in many state-of-the-art generative models and vision-language pipelines. In particular, CLIP’s ability to align visual and textual modalities has been widely adopted as a guiding mechanism in tasks ranging from text-to-image synthesis (e.g., GLIDE (Nichol et al., 2022), DALL·E (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022; Esser et al., 2024)) to creative applications such as style transfer and visual storytelling.
@@
Yet, despite its ubiquity, the nature and limits of CLIP’s perceptual alignment remain underexplored. While CLIP is optimized to match images with descriptive captions, it is less clear whether this alignment genuinely reflects human perception — especially in complex, subjective, or culturally embedded domains like art. Its capacity to identify what is depicted in an image is impressive, but its understanding of how that content is rendered — including factors such as artistic style, visual coherence, or historical context — is far less understood. Moreover, most evaluations have focused on natural images or benchmark datasets, leaving open the question of how well CLIP performs on ambiguous, interpretive, or aesthetic content, such as paintings — particularly those generated by AI.
@@
In this paper, we investigate CLIP’s perceptual capabilities in the domain of visual art, focusing on both human-made and AI-generated paintings. Our goal is to assess how well CLIP captures not only semantic content but also stylistic attributes, temporal signals, and visual deformations. Through a series of probing tasks and analyses, we evaluate CLIP’s representations along multiple interpretive axes, including scene type, artistic style, historical period, and the presence of visual artifacts. To this end, we leverage two richly annotated datasets: classical works from the National Gallery of Art of Washington (of Art, 2024) and synthetic paintings from the AI-Pastiche (Asperti et al., 2025a) collection.
@@
By combining metadata-driven evaluation with perceptual baselines, our work presents a structured critique of CLIP’s performance in aesthetic domains. In doing so, we highlight the need for more interpretable and perceptually grounded multimodal systems, particularly as models like CLIP are increasingly used to guide, evaluate, and curate creative content. As these systems become embedded in artistic workflows and cultural interfaces, it becomes crucial to ask not just what they can recognize or generate but how they see — and whether that way of seeing aligns with our own.
2. Related Works
Several recent works have focused on the problem of ascertaining the potential and limitations of CLIP’s vision mechanism and the extent to which the image embedding can be retrieved and effectively exploited through textual prompts.
Concerns about CLIP’s capacity to deeply understand the meaning of its textual prompts were raised in (Pan et al., 2022). In (Abbasi et al., 2025), a comprehensive analysis of CLIP’s performance limitations in multi-object contexts through controlled experiments is presented. The findings reveal significant biases in both encoders: the image encoder favors larger objects, while the text encoder prioritizes objects mentioned first in descriptions.
Several works have been devoted to assess the robustness of CLIP. In (Rosa et al., 2024), the authors investigate the adversarial robustness of CLIP-based methods in detecting AI-generated images, revealing their vulnerability to white-box attacks. A more comprehensive assessment of CLIP’s robustness along different perspectives is done in (Tu et al., 2024): despite several limitations, CLIP models exhibit superior robustness to visual factor-level variations compared to more traditional ImageNet models.
In (Lytvyn et al., 2024), the authors present a CLIP-based method for image similarity evaluation that relies on textual descriptions rather than raw visual features. Using cosine similarity between encoded text and image vectors, the approach captures semantic relationships more effectively than traditional metrics, such as SSIM and FSIM, especially in complex images. The results confirm CLIP’s robustness in tasks such as semantic search and classification, despite some limitations related to model dependence and real-world performance.
@@ 2. Related Works
A different line of research aimed to investigate the use of CLIP as a guidance technique for image generation. The findings in (Taghipour et al., 2024) suggest that while CLIP embeddings are important for aesthetic quality, they do not contribute significantly to the consistency of the subject and background in the generated outputs. Similar limitations in using CLIP encodings as a synthesis objective are discussed in (Liu et al., 2023), emphasizing the need for a more unified framework for text-guided and image-guided synthesis. None of these articles addresses guidance towards style replication; the conclusion of our work seems to suggest that CLIP is likely of little help for this task, if not altogether detrimental.
The possibility of understanding CLIP internal representations by inverting them with traditional gradient ascent techniques (Mahendran and Vedaldi, 2015; Dosovitskiy and Brox, 2016; Yin et al., 2020) has been explored in (Kazemi et al., 2024). Among other things, the authors remark on the frequent emergence of instances of Not Safe For Work (NSFW) images from the inversion of innocuous prompts. A gradient ascent technique, not based on pixel optimization but on a set of RGBA Bezier curves, is investigated in (Frans et al., 2022).
@@
In light of the previous limitations, several works have been devoted to improving CLIP’s performance, enhancing its downstream generalization ability, and reducing the modality gap between text and images. Adapters technique allow for a lightweight fine-tuning of the model through the insertions of suitable modules. Examples along this direction are CLIP-adapter (Gao et al., 2024), TIP-adapter (Yu et al., 2023), LIxP (Roth et al., 2024) or APE (Zhu et al., 2023). Given that our aim is to explore CLIP’s native perceptual abilities in the context of art, we consider adaptation techniques to be somewhat misaligned with the spirit of our investigation. A more promising line of investigation consists of addressing potential perception issues by focusing on subspaces through suitable projections, as outlined in (Zhu et al., 2024). However, our preliminary investigations in this direction did not yield interesting results.
@@
A different research direction consists of trying to improve the retrieval of multimodal systems through prompt engineering. A common approach followed e.g. in CoOp (Zhou et al., 2022b) or CoCoOp (Zhou et al., 2022a) consists of integrating the tokenization of the prompt with a set of learnable vectors learned by gradient descent. We tested this technique in the case of style classification, without noticeable improvements.
3. Methodology and Data
@@
We conducted our analyses on two richly annotated datasets: a subset of the freely available National Gallery of Art collection in Washington, and the AI-Pastiche dataset (Asperti et al., 2025a)—a collection of 953 AI-generated paintings produced by 12 different models from 73 carefully crafted prompts, covering a wide range of major artistic styles. Both datasets are described below.
The motivation for comparing human and AI-generated artworks lies in our goal to assess whether CLIP’s vision system can perceive the nuanced differences between them, including the limitations of AI models in accurately replicating artistic styles.
For the paintings from the National Gallery of Art, we leverage the available metadata to evaluate CLIP’s ability to associate each work with a descriptive summary and to understand its artistic style. We employ the Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes and Healy, 2018) as a visualization tool for exploring the shared image-text latent space of CLIP. An example of the resulting visualization, where the embeddings of the paintings are color-coded by their respective styles, is shown in Figure 1.
A similar investigation is conducted on the AI-Pastiche dataset, comparing the generated images both to their textual prompts and to the intended artistic styles. This analysis is inherently more complex, as any mismatch between image and prompt could result from either of the two models involved: the image generator, which may fail to accurately follow the prompt, or CLIP, which may fail to correctly identify the intended content or style. To disentangle these factors, we draw upon a set of user surveys conducted by the creators of the AI-Pastiche dataset. These surveys assess the perceived ”authenticity” of the AI-generated artworks, their adherence to the prompt, and the presence of visible artifacts or deformations.
Specifically, we use CLIP to evaluate how closely each generated image aligns with its corresponding prompt and compare these results with human evaluations. Additionally, we test CLIP’s capacity to identify visual defects, such as distortions, inconsistencies, or artifacts commonly produced by generative models.
@@
3.1. National Gallery of Art Dataset
The National Gallery of Art Dataset (NGAD) was created by carefully selecting 1,521 artworks from the National Gallery’s publicly available collection, which contains over 130,000 pieces(of Art, 2024). This selection focuses exclusively on paintings and drawings, deliberately excluding sculptures and modern artworks that might be more challenging for models like CLIP to analyze.
Each record in the dataset comprises 11 key attributes, including a unique identifier (objectid), title, creation period (period), artist or attribution, and a high-resolution image link (link) provided via the IIIF protocol. A key feature of the dataset is the description attribute, which provides a detailed textual representation of each artwork. These descriptions are structured to convey visual elements, artistic techniques, and compositional aspects.
The dataset covers a broad range of historical periods and artistic styles. The most represented styles include Baroque (286 artworks), Renaissance (272), and Impressionism (172), reflecting a focus on European artistic traditions.
While key attributes such as objectid, title, period, artist, link, and description are fully populated, certain fields contain missing values. Specifically, keyword is missing in 231 instances, style in 261 instances, and technique in 99 instances.
3.2. AI-Pastiche
The AI-Pastiche dataset (Asperti et al., 2025a) is a curated collection of 953 AI-generated paintings designed to mimic historical artistic styles. Created using 73 meticulously crafted prompts, the dataset spans a wide range of art periods. The images were generated using multiple state-of-the-art generative models and are intended to support the evaluation of how convincingly AI can reproduce the visual characteristics of human art. Each image is accompanied by rich metadata, including the generative model used, the prompt text, intended style and period, as well as a list of subject descriptors like ”crowd”, ”landscape” or ”soft tones.” The dataset enables research on stylistic imitation, model benchmarking, and user perception studies. It serves as a valuable resource for analyzing the capabilities and limitations of generative models in the artistic domain and is publicly available to support further exploration in AI-driven art.
4. Experiments on human artworks
The experiments in this section were conducted on a subset of the National Gallery of Art (Washington). We designed two distinct experiments. The first aims to assess CLIP’s ability to associate each image with a summary of its corresponding description. The second focuses on evaluating CLIP’s capacity to distinguish between different artistic styles.
4.1. Image-description alignment
In the first experiment, we encountered a limitation: CLIP accepts a maximum of 77 tokens as input, whereas the painting descriptions from the National Gallery of Art often exceeded this limit. To address this, we generated concise summaries of each description using ChatGPT 4o-mini. We passed as input to ChaptGPT also the subject, the style and the period of the painting, asking him to retain this information in the summary. The actual prompt used to generate the image is discussed in the appendix.
For each image-summary pair $\langle x,s\rangle$, we computed the cosine similarity between $CLIP_{image}(x)$ and $CLIP_{text}(s)$ and used this value as a ranking score to compute recall at different thresholds.
The experiments were repeated for several different versions of CLIP available in the OpenAI library(Radford et al., 2021).
The results of this experiment are shown in Table 1.
@@ 4.1. Image-description alignment
Table 1. Summary-image alignment for NGAD images.
These results show a consistent improvement in performance as the capacity of the CLIP models increases. Among the ResNet-based architectures, recall metrics progressively increase from RN50 up to RN50x64, with RN50x64 achieving the best performance in this family. Similarly, for the ViT-based models, the larger architectures and higher input resolutions result in better alignment, with ViT-L/14@336px showing the highest performance among the transformer variants.
Overall, while the ResNet-based RN50x64 yields the best recall@1 scores in this experiment, ViT-L/14@336px performs comparably, especially at higher recall thresholds. These results suggest that both architectural complexity and input resolution play a crucial role in enhancing the image-text alignment capabilities of CLIP, particularly in tasks involving fine-grained associations, such as matching painting summaries with artworks.
4.2. Style Recognition
In this second experiment, we assess CLIP’s ability to associate artworks with their corresponding artistic styles. For each unique style present in the dataset, we generated a fixed textual prompt in the form “an artwork in [style] style”. Using CLIP’s image and text encoders, we computed normalized embeddings for both modalities. The cosine similarity between text and image embeddings was then calculated to produce a similarity matrix, capturing the degree of alignment between each image and every style prompt.
Performance was evaluated using recall@$k$ metrics, which measure the proportion of test images for which the correct style appears among the top-$k$ most similar text prompts. Specifically, recall@$k$ is defined as the percentage of images for which the ground-truth style is ranked within the top $k$ most similar entries in the corresponding column of the similarity matrix.
Table 2 reports the results obtained across a range of CLIP architectures. Among all tested models, ViT-L/14@336px achieved the best performance, with a recall@1 of 0.457 and a recall@5 of 0.831. However, the results across all models remain moderate, highlighting the increased complexity of style recognition compared to textual description alignment.
@@ 4.2. Style Recognition
Table 2. Recall@k scores for art style recognition on NGAD.
To better understand the nature of the model’s errors, we present in Figure 2 the normalized confusion matrix obtained from the best-performing model. While the overall recall scores may appear modest, a closer inspection of the confusion matrix provides important nuance. Notably, many misclassifications occur between adjacent or stylistically related categories. For instance, Impressionism is frequently confused with Post-Impressionism, Abstract Expressionism with Minimalism, and Baroque with Rococo.
Interestingly, the Academic style appears to serve as a kind of fallback category, absorbing a diverse set of misclassifications from styles such as Baroque, Neoclassicism, Romanticism, and Realism. This suggests that the model may associate ”academic” with a broad range of classical or representational visual features, especially when more specific stylistic signals are ambiguous or absent.
Figure 2. Normalized confusion matrix of the best CLIP model (ViT-L/14@336px) on style classification.
This can be better observed in Figure 3, where we compare the distribution of true and predicted styles across the dataset. This analysis highlights notable disparities, reflecting both the imbalance present in the dataset and the most frequent mistakes in the model’s predictions.
Figure 3. Comparison between actual and predicted distribution of styles in NGAD.
@@
True: Impressionist Pred: Renaissance True: Realist Pred: Rococo True: Orientalist Pred: Neoclassic
Figure 4. Examples of misclassifications in style recognition
These examples illustrate how overlapping visual features or contextual ambiguity can lead to misclassifications. In several cases, the predicted style may appear visually plausible, emphasizing the challenge of the task.
4.3. Semantic Relationships in the Latent Space
To investigate the internal organization of the latent space learned by CLIP, we employed the UMAP algorithm (McInnes and Healy, 2018) to generate a three-dimensional projection of the high-dimensional embeddings. These embeddings included both image representations and textual prompts describing artistic styles (e.g., ”an artwork in [style] style”). The resulting visualization (Figure  5) reveals a clear separation between textual encodings (the small cluster on the left) and image encodings (the large cluster on the right). For the images, we also distinguish correctly classified samples, shown as green bullets, from misclassified ones, shown as red crosses.

### v2

@@
Abstract
Multimodal systems and Large Language Models have shown remarkable capabilities in text-based reasoning, yet their capacity to perceive and interpret visual art remains uncertain. This study examines how CLIP “sees” and understands artworks by comparing their responses to human- and AI-generated paintings in the European tradition from the Renaissance onward. The analysis focuses on its ability to identify style, period and cultural context, as well as potential biases in its perception, evaluated against human judgments.
@@ Abstract
Figure 1: 3D UMAP projection of image embeddings extracted by the CLIP ViT-L/14 model from the National Gallery of Art.
I Introduction
@@
In recent years, multimodal models have reshaped the landscape of machine perception and understanding, with CLIP (Contrastive Language–Image Pretraining) [23] standing out as one of the most influential. Trained on hundreds of millions of image–text pairs, CLIP has demonstrated remarkable capabilities across a broad range of tasks, including zero-shot classification [31, 12, 33], image retrieval [32, 5, 15], re-identification [13, 4], prompt-based generation [28, 10], and semantic search [16, 22, 38]. Its success has made it a foundational component in many state-of-the-art generative models and vision-language pipelines. In particular, CLIP’s ability to align visual and textual modalities has been widely adopted as a guiding mechanism in tasks ranging from text-to-image synthesis (e.g., GLIDE [19], DALL·E [24], Stable Diffusion [25, 7]) to creative applications such as style transfer and visual storytelling.
@@
Yet, despite its ubiquity, the nature and limits of CLIP’s perceptual alignment remain underexplored. While CLIP is optimized to match images with descriptive captions, it is less clear whether this alignment genuinely reflects human perception — especially in complex, subjective, or culturally embedded domains like art. Its capacity to identify what is depicted in an image is remarkable, but its understanding of how that content is rendered — including factors such as artistic style, visual coherence, or historical context — is far less understood.
@@
In this paper, we investigate CLIP’s perceptual capabilities in the domain of visual art, focusing on both human-made and AI-generated paintings. We treat CLIP’s vision encoder as a fixed perceptual system, without any fine-tuning or modification, in analogy to the human sensory apparatus. Our goal is to assess how well CLIP captures not only semantic content but also stylistic attributes, temporal signals, and visual deformations. Through a series of probing tasks and analyses, we evaluate CLIP’s representations along multiple interpretive axes, including scene type, artistic style, historical period, and the presence of visual artifacts. To this end, we leverage two richly annotated datasets: classical works from the National Gallery of Art of Washington [20] and synthetic paintings from the AI-Pastiche [3] collection.
@@
By combining metadata-driven evaluation with perceptual baselines, our work presents a structured critique of CLIP’s performance in aesthetic domains. In doing so, we highlight the need for more interpretable and perceptually grounded multimodal systems, particularly as models like CLIP are increasingly used to guide, evaluate, and curate creative content.
II Related Works
Several recent studies have examined the potential and limitations of CLIP’s vision mechanism, particularly how well its image embeddings can be retrieved and exploited via textual prompts. Concerns about CLIP’s ability to fully grasp prompt semantics were raised in [21]. In [1], controlled experiments in multi-object contexts revealed notable biases: the image encoder favors larger objects, while the text encoder prioritizes objects mentioned first. CLIP’s robustness has also been investigated. [26] showed that CLIP-based methods for detecting AI-generated images are vulnerable to white-box attacks, while [30] offered a broader evaluation, finding that despite limitations, CLIP models are generally more robust to visual factor variations than traditional ImageNet models.
@@ II Related Works
Another line of research has investigated the use of CLIP as a guidance technique for image generation. Findings in [29] indicate that while CLIP embeddings contribute to aesthetic quality, they have limited influence on maintaining consistency between subject and background. Comparable limitations are discussed in [14], which emphasizes the need for a more unified framework integrating text- and image-guided synthesis. However, none of these studies considers guidance for style replication, and the results presented here suggest that CLIP offers limited benefit for this purpose and may, in some cases, be counterproductive.
The possibility of understanding CLIP internal representations by inverting them with traditional gradient ascent techniques [17, 6, 34] has been explored in [11]. A gradient ascent technique, not based on pixel optimization but on a set of RGBA Bezier curves, is investigated in [8]. All these studies seem to highlight the fact that only basic semantic information is encoded in CLIP’s embeddings.
@@
In light of the previous limitations, several works have been devoted to improving CLIP’s performance, enhancing its downstream generalization ability, and reducing the modality gap between text and images. Adapters technique allows for a lightweight fine-tuning of the model through the insertion of suitable modules. Examples along this direction are CLIP-adapter [9], TIP-adapter [35], LIxP [27] or APE [39]. Given that our aim is to explore CLIP’s native perceptual abilities in the context of art, we consider adaptation techniques to be somewhat misaligned with the spirit of our investigation. A line of investigation closer to our objectives consists of addressing potential perception issues by focusing on subspaces through suitable projections, as outlined in [40]. However, our preliminary investigations in this direction did not yield interesting results.
@@
A different research direction consists of improving the discrimination and retrieval capacities of multimodal systems through prompt engineering. A common followed approach, e.g. in CoOp [37] or CoCoOp [36], consists of integrating the tokenization of the prompt with a set of learnable vectors learned by gradient descent. We tested this technique in the case of style classification, without noticeable improvements.
III Methodology and Data
@@
We conducted our analyses on two richly annotated datasets: a subset of the freely available National Gallery of Art collection in Washington, and the AI-Pastiche dataset [3]—a collection of 953 AI-generated paintings produced by 12 different models from 73 carefully crafted prompts, covering a wide range of major artistic styles. The motivation for comparing human and AI-generated artworks lies in our goal to assess whether CLIP’s vision system can perceive the nuanced differences between them, including the limitations of AI models in accurately replicating artistic styles.
For the paintings from the National Gallery of Art, we leverage the available metadata to evaluate CLIP’s ability to associate each work with a descriptive summary and to understand its artistic style. We employ the Uniform Manifold Approximation and Projection (UMAP) algorithm [18] as a visualization tool for exploring the shared image-text latent space of CLIP. An example of the resulting visualization, where the embeddings of the paintings are color-coded by their respective styles, is shown in Figure 1.
A similar investigation is conducted on the AI-Pastiche dataset, comparing the generated images to both their textual prompts and to the intended artistic styles. This analysis is inherently more complex, as any mismatch between image and prompt could result from either of the two models involved: the image generator, which may fail to accurately follow the prompt, or CLIP, which may fail to correctly identify the intended content or style. To unravel these factors, we draw on a set of user surveys conducted by the creators of the AI-Pastiche dataset. These surveys assess the perceived ”authenticity” of AI-generated artworks, their adherence to the prompt, and the presence of visible artifacts or deformations.
Specifically, we use CLIP to evaluate how closely each generated image aligns with its corresponding prompt and compare these results with human evaluations. In addition, we test CLIP’s capacity to identify visual defects, such as distortions, inconsistencies, or artifacts commonly produced by generative models.
@@
IV Experiments on human artworks
The experiments in this section were conducted on a subset of the National Gallery of Art in Washington. We designed two distinct experiments. The first aims to assess CLIP’s ability to associate each image with a summary of its corresponding description. The second focuses on evaluating CLIP’s capacity to distinguish between different artistic styles.
IV-A Image-description alignment
In the first experiment, a limitation arose from CLIP’s 77-token input cap, as painting descriptions from the National Gallery of Art often exceeded this length. To address this, concise summaries (x) were generated using ChatGPT-4o-mini, ensuring that content, style, and period information from the original description was preserved. The actual image-generation prompt is provided in the extended version [2]. For each image–summary pair ⟨x, s⟩, we computed the cosine similarity between $CLIP_{image}(x)$ and $CLIP_{text}(s)$ and used this value as a ranking score to compute recall at different thresholds. The experiments were repeated for several different versions of CLIP available in the OpenAI library[23]; results are reported in Table I.
@@ IV-A Image-description alignment
TABLE I: Summary-image alignment for NGAD images.
The results show a consistent improvement in performance as the capacity of the CLIP models increases. Overall, while the ResNet-based RN50x64 yields the best recall@1 scores in this experiment, ViT-L/14@336px performs comparably, especially at higher recall thresholds. This suggests that both architectural complexity and input resolution play a crucial role in enhancing the image-text alignment capabilities of CLIP, particularly in tasks involving fine-grained associations, such as matching painting summaries with artworks.
IV-B Style Recognition
In this second experiment, we assess CLIP’s ability to associate artworks with their corresponding artistic styles. For each unique style present in the dataset, we generated a fixed textual prompt in the form “an artwork in [style] style”. Using CLIP’s image and text encoders, we computed normalized embeddings for both modalities. The cosine similarity between text and image embeddings was then calculated to produce a similarity matrix, capturing the degree of alignment between each image and every style prompt, at different recall thresholds.
Table II reports the results obtained across a range of CLIP architectures. Among all tested models, ViT-L/14@336px achieved the best performance, with a recall@1 of 0.457 and a recall@5 of 0.831. However, the results across all models remain moderate, highlighting the increased complexity of style recognition compared to textual description alignment.
@@ IV-B Style Recognition
TABLE II: Style recognition scores on NGAD.
In Figure 2, we compare the distribution of true and predicted styles across the dataset. This analysis highlights notable disparities, reflecting both the imbalance present in the dataset and the most frequent mistakes in the model’s predictions.
Figure 2: Comparison between actual and predicted distribution of styles in NGAD.
@@
True: Impressionist
Pred: Renaissance
True: Realist
Pred: Rococo
True: Orientalist
Pred: Neoclassic
Figure 3: Examples of misclassifications in style recognition
These examples illustrate how overlapping visual features or contextual ambiguity can lead to misclassifications.
IV-C Semantic Relationships in the Latent Space
To investigate the internal organization of the latent space learned by CLIP, we employed the UMAP algorithm [18] to generate a three-dimensional projection of the high-dimensional embeddings. These embeddings included both image representations and textual prompts describing artistic styles (e.g., ”an artwork in [style] style”). The resulting visualization (Figure  4) reveals a clear separation between textual encodings (the small cluster on the left) and image encodings (the large cluster on the right). For the images, we also distinguish correctly classified samples, shown as green bullets, from misclassified ones, shown as red crosses.